Department B had a beginning inventory of 2,200 units 1/4 completed an ending inventory of 2,100 units 2/3 completed, and received 1,500 units during the period from department A. What were the equivalent units of production of Department B, assuming weighted-average process costing?
Answer #1

Units completed = Beginning inventory + Units received - Ending inventory

= 2,200 + 1,500 - 2,100

= 1,600

Equivalent ending units = 2,100 X 2/3 = 1,400

Equivalent units = Units completed + Equivalent ending units

= 1,600 + 1,400

= 3,000